OpenStudy (anonymous):

Write the given expression as an algebraic expression in x. cos(2 tan−1 x)

OpenStudy (anonymous):

Use the Half angle and Double angle properties to write in terms of x.

OpenStudy (tkhunny):

Yes, that would do it. \(\cos(2x) = 2\cos^{2}(x) - 1 = 1 - 2\sin^{2}(x)\)